OpenStudy (natalieee):

anybody have geometry 10a ? i need help with lesson 6 quiz: trapezoids and kites.

OpenStudy (anonymous):

i took this test today

OpenStudy (natalieee):

could you help me please ?

OpenStudy (anonymous):

yes i just need 10 but ik the rest

OpenStudy (natalieee):

i need 7 8 9 10.

OpenStudy (anonymous):

7. a=144 b = 67 8. lm=85.5 9.m<1= 39 m<3=51

OpenStudy (natalieee):

thanks you so much! i will let you know if i get number 10.

OpenStudy (anonymous):

np and ok thanks

OpenStudy (natalieee):

@stewie hey you there ? how did you get 85.5 for 8 ?

OpenStudy (anonymous):

The length of the median is the average length of the bases so then LM = (AB + DC) / 2 LM = (46 + 125) / 2 LM = 85.5

OpenStudy (natalieee):

that makes sense. thank you.

OpenStudy (anonymous):

10 is 16

OpenStudy (anonymous):

np

OpenStudy (natalieee):

(:

OpenStudy (anonymous):

10 is 12 i mean sorry

OpenStudy (natalieee):

okay, how did you get 12 ?

OpenStudy (anonymous):

x=7 and y= 16 you can also add this but if you plug them in for x n y you should get 12 for everyone